# Ice Cream Doesn’t Cause Drowning: Benchmarking LLMs Against Statistical Pitfalls in Causal Inference

### v1

@@ 4 Results and Analysis
In this section, we present the key findings from our experiments, examining the causal inference capabilities of five state-of-the-art LLMs (GPT-4o, GPT-o4-mini [7], claude-3.5-sonnet [38], gemini-2.0 flash [39], and Deepseek-chat [40]). We examine these LLMs using two protocols: direct prompting and code-assisted prompting, tested on 6 categories of pitfalls and corresponding 15 challenges, each with 4 datasets and 5 questions of varying difficulty levels.
Overall Performance.
Our evaluation revealed that GPT-o4-mini demonstrated robust performance overall, achieving the highest average causal reliability scores across the majority of pitfalls. Specifically, it obtained an average of $40.02\%$ in direct prompting and improved to $44.63\%$ with code-assisted prompting (Table 5, Figure 1). Meanwhile, in the pitfall category interventions and experimental reasoning, Deepseek-chat achieved superior performance ($52.42\%$) under direct prompting compared to GPT-o4-mini ($45.21\%$).
@@
Claude-3.5-sonnet 14.82 47.60 12.00 40.50 14.94 19.47 24.89
Gemini-2.0-flash 16.57 37.57 13.43 46.67 10.94 10.93 22.68
Deepseek-chat 15.20 52.42 12.86 53.83 20.83 29.67 30.80
GPT-4o 24.71 34.86 12.57 33.50 12.48 11.73 21.64
GPT-o4-mini 37.33 45.21 18.57 57.67 38.12 43.20 40.02
@@
Claude-3.5-sonnet 20.41 45.18 11.71 41.96 14.29 22.64 26.03
Gemini-2.0-flash 30.57 42.71 14.29 42.17 15.18 39.73 30.77
Deepseek-chat 18.40 47.25 10.86 47.13 25.79 51.40 33.47
GPT-4o 23.14 40.17 13.14 36.33 19.10 48.80 30.11
GPT-o4-mini 61.67 52.59 17.33 50.36 25.06 60.78 44.63
@@
Introducing computational assistance through code-assisted prompting enhanced LLM reliability in most categories. For example, in the categories of confounding bias and external validity, GPT-o4-mini’s performance significantly improved from $37.33\%$ and $61.67\%$ in direct prompting to $43.20\%$ and $60.78\%$, respectively, with code-assisted prompting (Table 5). Similar gains were observed across other models, highlighting the significant benefit of incorporating computational procedures when addressing complex causal tasks.
@@
Analyzing LLM performance across difficulty levels (Table 6), we observed that code-assisted prompting was especially beneficial for harder questions. GPT-o4-mini, for example, improved substantially from $17.84\%$ (direct prompting) to $30.26\%$ (code-assisted prompting) on very hard tasks. This improvement illustrates the value of computational guidance when navigating challenging causal scenarios. However, even with such assistance, the absolute scores on the hardest tasks remain relatively low, indicating the complexity of reliable causal reasoning.
Table 6: Causal reliability by difficulty levels of questions, comparing direct and code-assisted prompting. Values represent averages of normalized score, defined in equation (1), across 16 challenges; higher scores indicate better performance.
@@
Claude-3.5-sonnet 39.37 36.14 31.27 15.11 8.81
Gemini-2.0-flash 37.63 32.76 27.53 16.82 8.28
Deepseek-chat 45.60 37.77 37.55 29.87 16.52
GPT-4o 33.72 29.28 24.44 12.16 9.47
GPT-o4-mini 59.93 53.97 49.24 30.74 17.84
@@
Claude-3.5-sonnet 36.82 35.78 31.93 20.12 17.08
Gemini-2.0-flash 46.47 39.81 31.80 24.50 21.64
Deepseek-chat 52.94 47.77 42.90 26.11 14.29
GPT-4o 49.52 41.64 37.18 21.91 11.65
GPT-o4-mini 55.31 48.80 48.36 30.78 30.26
@@
Overall, our findings emphasize significant reliability gaps persisting across all evaluated models and pitfalls. Even the best-performing model, GPT-o4-mini, achieved a modest $44.63\%$ average score under code-assisted prompting (Table 5). These consistent gaps indicate substantial opportunities for future improvements.
Human-LLM grading alignment
@@
As shown in Figure 4, code-execution failures peak in the “mediation effects” and “external validity” categories, where implementing correct stratification and transportability routines is most demanding. Interestingly, “very easy” questions produce the highest failure rates, whereas “very hard” questions yield lower rates. One possible reason is that the easiest questions often include detailed hints and expert knowledge that encourage the model to generate more complex, specific code. Consequently, this complexity increases the chance of syntax or logic errors, indicating the need for more robust code generation and automated error-handling in future code-assisted prompting frameworks.

### v2

@@ 4 Results and Analysis
In this section, we present the key findings from our experiments, examining the causal inference capabilities of ten large language models. Our evaluation spans both closed source and open source LLMs, with the complete list of models provided in the Appendix. All models were evaluated using two protocols: direct prompting and code assisted prompting. The assessment covered six categories of causal inference pitfalls and fifteen challenges, each built from five datasets and five questions of varying difficulty.
@@
Gemma2-9b 14.00 30.72 8.00 15.89 9.49 4.05 13.69
Llama3.1-8b 18.46 34.78 10.86 29.33 7.71 6.05 17.86
Llama3.1-70b 17.60 35.92 10.29 33.67 11.31 9.92 19.78
Mistral-7b 17.31 29.83 5.71 19.17 8.40 6.18 14.43
Mixtral-8x22b 16.40 32.08 8.00 27.50 9.20 8.75 16.99
Claude-3.5-sonnet 18.74 47.60 12.00 40.50 18.63 19.84 26.22
Gemini-2.0-flash 20.06 37.56 13.43 46.67 13.48 14.93 24.36
Deepseek-chat 25.89 52.42 12.86 53.83 20.83 28.74 32.43
GPT-4.1 17.26 33.57 6.57 53.27 16.43 24.31 25.24
GPT-o4-mini 41.43 45.21 18.57 57.67 36.97 44.48 40.72
@@
Gemma2-9b 7.89 18.50 4.86 23.04 6.63 18.50 13.24
Llama3.1-8b 11.93 15.99 7.67 18.06 6.23 17.29 12.86
Llama3.1-70b 23.54 22.83 8.21 25.17 14.41 25.17 19.89
Mistral-7b 4.68 13.15 1.43 11.11 6.43 9.09 7.65
Mixtral-8x22b 22.50 30.53 5.44 26.49 14.29 24.31 20.59
Claude-3.5-sonnet 32.91 47.33 11.71 41.73 16.12 26.47 29.38
Gemini-2.0-flash 37.20 42.96 14.29 42.17 16.19 37.98 31.80
Deepseek-chat 38.63 48.70 10.86 47.13 25.79 45.63 36.12
GPT-4.1 47.14 42.65 12.29 49.40 23.87 48.58 37.32
GPT-o4-mini 62.00 51.86 16.96 50.00 26.67 50.71 43.03
@@
Overall Performance.
Across models, GPT-o4-mini demonstrated the highest overall reliability, achieving an average causal reliability of $40.72\%$ under direct prompting and $43.03\%$ under code-assisted prompting (Table 4). Deepseek-chat, although lower in aggregate (average 32.43% under direct prompting and 36.12% under code-assisted prompting), obtained the strongest performance in interventions and experimental reasoning (52.42% under direct prompting and 48.70% under code-assisted prompting). Taken together, these results indicate that optimized mid-scale models can, in certain contexts, outperform larger frontier systems on causal reasoning tasks.
@@
Although code-assisted prompting improved several strong models, its benefits were not universal. For example, GPT-o4-mini and Gemini-2.0-flash gained from code execution, increasing their averages from $40.72\%$ to $43.03\%$ and $24.36\%$ to $31.80\%$, respectively. GPT-4.1 also improved significantly ($25.24\%\rightarrow 37.32\%$). Under code-assisted prompting, models first convert raw data into summary statistics through generated code, then reason over those statistics. This separates low-level data parsing from causal reasoning, which partly explains why stronger models benefit more: they produce correct analysis code and reason on clean numerical outputs rather than raw tables. In contrast, some small open-source LLMs showed little benefit: Llama3.1-8B decreased from $17.86\%$ to $12.86\%$, Gemma2-9B remained unchanged ($13.69\%$ to $13.24\%$), and Mistral-7B even dropped from $14.43\%$ to $7.65\%$. These results indicate that computational results increase the strengths of already-capable LLMs, but does not provide a uniform advantage across the LLMs. Models with high code-error rates (e.g., Mistral-7B, Llama-8B) have worse scores under code-assisted prompting. Allowing one debugging attempt can bring their performance back to direct-prompting levels. We discuss this further in Appendix H.
@@
Causal reliability consistently decreases as the question becomes harder. As shown in Table 5, all models performed best on very easy and easy questions, with average scores declining as task difficulty increased, where the question contains fewer hints (see Table 2. For example, GPT-o4-mini achieved $60.72\%$ under direct prompting and $56.73\%$ under code-assisted prompting on very easy items, but dropped to $17.75\%$ and $32.84\%$, respectively, on very hard questions. Deepseek-chat and Gemini-2.0-flash showed similar trends, maintaining moderate performance on medium and hard levels but falling quickly on the very hard questions. These results indicate that code assistance is particularly beneficial for challenging tasks, helping stronger models recover some performance at higher difficulty levels.
Table 5: Causal reliability by difficulty levels of questions, comparing direct and code-assisted prompting. Values represent averages of normalized score, defined in equation (1), across 15 challenges; higher scores indicate better performance.
@@
Gemma2-9b 20.50 15.64 14.22 6.72 5.46
Llama3.1-8b 27.99 20.85 19.90 10.71 5.82
Llama3.1-70b 31.04 27.38 22.04 11.29 5.23
Mistral-7b 20.04 18.88 18.02 7.49 3.80
Mixtral-8x22b 27.80 21.20 20.31 8.41 5.23
Claude-3.5-sonnet 40.27 36.63 29.69 13.42 8.30
Gemini-2.0-flash 37.83 33.59 27.59 15.72 8.22
Deepseek-chat 46.37 38.92 37.92 28.79 16.36
GPT-4.1 38.01 33.94 33.59 17.31 7.67
GPT-o4-mini 60.72 54.05 48.13 30.72 17.75
@@
Gemma2-9b 19.09 18.35 16.43 10.56 7.94
Llama3.1-8b 26.04 14.41 12.77 8.84 7.90
Llama3.1-70b 32.21 27.88 24.77 14.08 7.58
Mistral-7b 9.44 12.03 9.59 6.63 2.73
Mixtral-8x22b 33.14 29.77 25.64 14.59 7.77
Claude-3.5-sonnet 39.70 35.06 33.55 20.92 18.34
Gemini-2.0-flash 46.01 39.83 32.17 28.00 21.67
Deepseek-chat 55.56 48.23 43.69 28.46 16.50
GPT-4.1 57.80 46.49 46.48 24.33 19.65
GPT-o4-mini 56.73 49.66 49.37 34.31 32.84
@@
Despite variation across models and prompting strategies, significant reliability gaps remain in all settings. Among the tested LLMs, GPT-o4-mini achieved the highest causal reliability score $43.03\%$ on average under code-assisted prompting, while most other models remained well below this level. Performance on difficult causal reasoning challenges was especially limited, with very hard questions rarely exceeding $30\%$ even for the best LLMs. Taken together, the results suggest that existing LLMs, when used without finetuning or specialized architectures, remain unreliable to apply in high-stakes causal inference.
Human-LLM Grading Alignment.
@@
As shown in Figure 4, code‐execution failures peak in the “mediation effects” and “interventions and experimental reasoning” categories, where implementing correct stratification and transportability routines is most demanding. Interestingly, “very easy” questions produce the highest failure rates, whereas “very hard” questions yield lower rates.